The theoretical probability of undesirable side effects resulting from taking Pyrene is 1 in 15. If 180 people take Pyrene to relieve congestion, theoretically how many will encounter undesirable side effects?
Answer by edjones(8007)

x/180 = 1/15
x=180/15
=12
.
Ed